# Beyond Fidelity: Explaining Vulnerability Localization of Learning-based Detectors

### 05 Jan 2024 v1

@@ Abstract.
Abstract: Vulnerability detectors based on deep learning (DL) models have proven their effectiveness in recent years. However, the shroud of opacity surrounding the decision-making process of these detectors makes it difficult for security analysts to comprehend. To address this, various explanation approaches have been proposed to explain the predictions by highlighting important features, which have been demonstrated effective in other domains such as computer vision and natural language processing. Unfortunately, an in-depth evaluation of vulnerability-critical features, such as fine-grained vulnerability-related code lines, learned and understood by these explanation approaches remains lacking. In this study, we first evaluate the performance of ten explanation approaches for vulnerability detectors based on graph and sequence representations, measured by two quantitative metrics including fidelity and vulnerability line coverage rate. Our results show that fidelity alone is not sufficient for evaluating these approaches, as fidelity incurs significant fluctuations across different datasets and detectors. We subsequently check the precision of the vulnerability-related code lines reported by the explanation approaches, and find poor accuracy in this task among all of them. This can be attributed to the inefficiency of explainers in selecting important features and the presence of irrelevant artifacts learned by DL-based detectors.
@@ 1. Introduction
The ubiquity of software in today’s digital society underpins the paramount importance of vulnerability detection techniques. As we propel further into the digital age, software permeates every corner of our lives, from healthcare to finance, transportation to communication. This pervasive deployment has led to an exponential increase in the surface area for potential attacks, amplifying the demand for robust, accurate, and efficient vulnerability detection mechanisms. Current vulnerability detection techniques (Coverity, ; RATS, ; Infer, ; CSA, ; Fortify, ; Checkmarx, ; SVF, ; devign, ; reveal, ; ivdetect, ; dwk, ; SySeVR, ; vdp, ; vulcnn, ; Russel, ), however, are often characterized by significant limitations that stymie their effectiveness. A core shortcoming lies in their limited ability to pinpoint the exact code lines that instigate vulnerabilities. Existing detection tools frequently flag entire segments of code as potentially hazardous, leading to vast sections needing review. This wide net approach not only overwhelms developers with high rates of false positives but also sows the seeds of inefficiency and can potentially delay mitigation efforts. The ability to identify the precise lines of code responsible for a vulnerability, thus, is an issue of critical importance and poses an intricate challenge in the realm of program analysis.
@@
In our study, the overarching objective is to conduct a rigorous evaluation and exmaination of the performance of these explanation techniques within the field of fine-grained vulnerability localization. we set our sights on evaluating five graph-based explanation methodologies: GNNExplainer (gnnexplainer, ), PGExplainer (pgexplainer, ), GNN-LRP (gnnlrp, ), GradCAM (grad, ), and DeepLift (deeplift, ). These techniques are examined in conjunction with four graph-based vulnerability detectors, namely DeepWuKong (dwk, ), IVDetect (ivdetect, ), Reveal (reveal, ), and Devign (devign, ). Furthermore, we venture into the investigation of five sequence-based techniques—LRP (LRP-LSTM, ), Shap (SHAP, ), Lime (LIME, ), GradInput (GradInput, ), and DeepLift (deeplift, )—in relation to sequence-based vulnerability detectors such as VulDeePecker (vdp, ), SySeVR (SySeVR, ), and TokenLSTM.
@@
In our evaluation process, we initially use fidelity as the metric to evaluate existing explanation techniques, a common practice in explainable AI to assess the accuracy of generated explanations. However, we find that relying solely on fidelity leads to significant fluctuations in scores across different detectors and datasets, showing inconsistent results. To address this, we introduce an additional metric called Vulnerability Line Coverage rate (LC), mirroring MSR (ContraFlow, ), which divides into two sub-metrics: the Vulnerability-Triggering Code Line Coverage rate (TLC) and the Vulnerability-Fixing Code Line Coverage rate (FLC). These aim to provide a more detailed understanding of how well the explanation techniques can identify and explain vulnerabilities. However, our findings show that most explanation techniques struggle to pinpoint key vulnerability-related statements accurately, as seen in both TLC and FLC measurements (RQ2, Section 5.4). For instance, when applying GNNExplainer to IVDetect on the SARD dataset, the results show a wide range of TLC values from 7% to 64%, and FLC values from 18% to 66%, despite the fidelity scores remaining relatively high between 74% and 92%. This highlights that existing approaches, even with high fidelity scores, fail to learn and reflect the critical features inherent to the vulnerabilities, indicating that a broader evaluation beyond fidelity is necessary to truly understand the effectiveness of explanation techniques in addressing vulnerabilities.
@@
Throughout our study, we have gained two in-depth understandings regarding the DL-based vulnerability detectors and explanation approaches respectively. On the one hand, the existing explanation approaches are prone to pick up vulnerability-irrelevant features as explanations. On the other hand, deep learning-based detectors indeed perform vulnerability detection by diligently learning the distinctions between vulnerable and normal samples. The varying sensitivities of model-related explainers on retrained and previous models to vulnerability-fixing lines suggest that they may primarily identify feature discrepancies between positive and negative samples, rather than understanding how vulnerabilities are triggered (RQ3.1, Section 5.5.1). To further validate our point and evaluate their comprehension of vulnerabilities, we manually patch a subset of vulnerabilities, using the originally patched samples as a reference. By comparing detection accuracy between non-vulnerability samples, we anticipate minimal accuracy differences if the detectors truly understand vulnerability nuances (RQ 3.2, Section 5.5.2). Surprisingly, a significant portion of manually fixed codes is misclassified as vulnerable, with fewer false alarms in the originally patched samples. Moreover, we identify a kind of irrelevant text features that notably influence the model’s predictions, collectively indicating that current detectors have not fully understood vulnerability semantics.
@@
  • Enhanced Evaluation Metrics. We employ the quantitative metric of fidelity and introduce additional metrics, namely TLC and FLC, to evaluate the efficacy of the explanation approaches in accurately localizing vulnerabilities. Our investigation reveals that none of the explanation approaches effectively localize vulnerabilities even when they achieve high fidelity scores.
  • Revealing Limitations in Deep-Learning Vulnerability Detectors’ Capabilities. We find that deep-learning vulnerability detectors aim to identify vulnerabilities by learning distinctions between vulnerable and normal samples. Improved learning leads to heightened sensitivity to distinctions between black and white samples. Gradient-based and decomposition-based methods can capture this model-embedded information. However, models can still be misdirected by irrelevant disturbances, hinting at the inadequacy of existing perturbation-based explanation methods and evaluation metric fidelity for vulnerability detection scenarios.
  • Revealing Limitations in Recognizing Fixing Patterns and Text Feature Disturbances. Upon further assessment of the DL-based detectors, it becomes evident that these models are unable to recognize fixing patterns not present in the training dataset. Simultaneously, we identify a specific text feature that can trigger false positives in the model. Indicating DL-based detectors primarily distinguish between vulnerable and normal samples without a full comprehension of the mechanisms underlying vulnerabilities.
  • New Evaluation Metric for DL-based detectors. To complement the existing evaluation metrics, we introduce a metric namely VUR that focuses on the impact of vulnerability-unrelated perturbations on DL-based detectors. The metric is employed to determine the proportion of perturbations that are unrelated to vulnerabilities. We observe that all detectors are affected to varying extents by vulnerability-unrelated perturbations, causing a flip in their prediction outcomes.
@@ 2. Background
In this section, we mainly introduce the DL-based detectors and explanation approaches used in this study.
@@ 2.1. DL-based Vulnerability Detectors
In the realm of vulnerability detection, deep learning-based approaches have emerged as potent tools for identifying security vulnerabilities within software programs. These methodologies typically entail segmenting source code into discrete code fragments, subsequently vectorizing each fragment using deep learning models, and predicting the likelihood of a vulnerability existing within that fragment. The inherent advantage of these methods lies in their capacity to autonomously discern patterns and regularities within source code, thereby enhancing the accuracy and efficiency of vulnerability detection.
Based on the manner in which source code is segmented, deep learning-based vulnerability detectors can be categorized into two primary classes: function-level and slice-level detectors. In function-level detectors (Russel, ; funded, ; reveal, ; ivdetect, ; devign, ; linevd, ; linevul, ), individual functions within the program are treated as isolated code fragments. In contrast, slice-level detectors (vulspg, ; dwk, ; vdp, ; SySeVR, ; vuldeelocator, ) extract code fragments spanning multiple functions based on data dependencies or control flow dependencies of vulnerability-critical statements. This approach offers a comprehensive contextual view of vulnerabilities, thus contributing to improved detection precision.
Furthermore, with respect to the vectorization of code fragments, these vulnerability detectors can be further classified as sequence-based and graph-based detectors. Sequence-based methods directly parse source code fragments into token sequences and employ sequence neural networks for vectorization. Conversely, graph-based approaches typically parse the structures of code fragments, such as abstract syntax tree (AST), control flow graphs (CFG), or program dependency graphs (PDG), and further process these graph structures using graph neural networks. This facilitates the capture of structural information and dependency relationships within the source code, thereby enhancing the effectiveness of vulnerability detection.
@@
Table 1. Overview of seven learning-based detectors.
@@ 2.2. Explanation Approaches
  • Gradient-based (grad, ) explanation methods quantify the impact of a feature, denoted as $x_{i}$, in the input sample on the model’s output, $o$, by calculating the partial derivative of $o$ concerning feature $i$. In this research, we opt for $grad_{input}$ (GradInput, ), which represents a gradient-based method and is calculated as the dot product of the input vector and the partial gradient.
  • LRP (i.e. layer-wise relevance propagation) represents another significant approach for leveraging the internal relationships within neural networks to estimate feature importance scores. In essence, LRP utilizes the backpropagation mechanism of neural networks to propagate the output scores (relevance) from the model’s output layer all the way back to the input layer. Consequently, each input feature receives a relevance value, serving as an importance score. To apply the LRP method to LSTMs and GRUs, we employ the propagation rule introduced by Arras Leila et al (LRP-LSTM, ).
@@
Table 2. Basic information explainers
Explanation Method Representation Introduction GNNExplainer (gnnexplainer, ) Graph GNNExplainer trains mask for edges and node features to identify meaningful input information. These masks are updated to maximize the mutual information between original and updated predictions. PGExplainer (pgexplainer, ) Graph PGExplainer trains a parameterized mask predictor to predict the edge mask value. The predictor takes embedding vectors of two nodes connected by an edge as input and produces a corresponding mask value for the edge. GNN-LRP (gnnlrp, ) Graph GNN-LRP explores the importance of various graph walks through score decomposition. It decomposes model predictions into different graph walks using LRP GradCAM (grad, ) Graph Pope, Phillip E et al. (grad, ) extend GradCAM (gradcam, ) to the realm of graph classification, where it initially calculates the gradients of the target prediction with respect to the final node embeddings. Subsequently, these gradients are averaged to derive the weights for each feature map. DeepLift (deeplift, ) Both DeepLift combines LRP and Integrate Gradient, utilizing their insights to enhance its backpropagation process. By incorporating a reference input, DeepLift overcomes discontinuities and achieves superior performance in various tasks compared to traditional LRP methods. It is extended to support deep graph models by Yuan et al. (explain_survey, ). Lime (LIME, ) Sequence Lime employ a surrogate model to approximate the performance of the target model. By perturbing the input $x$, LIME samples neighboring points ${x_{1}^{{}^{\prime}},x_{2}^{{}^{\prime}},...,x_{n}^{{}^{\prime}}}$ and obtains their labels through the target model. Shap (SHAP, ) Sequence Shap estimates the Shapley values of each feature in the prediction by adjusting the loss function, weight kernel, and solving a regression problem, thus providing additive explanations. Gradient * Input (GradInput, ) Sequence Gradient * Input is an early-stage gradient-based attribution method that computes attribution by taking the partial derivative of the output with respect to the input and multiplying it by the input itself. LRP (LRP-LSTM, ) Sequence LRP is a technique that approximates feature importance scores by leveraging the internal relationship between input and output in neural networks. It uses backpropagation to propagate relevance scores from the output layer to the input layer, assigning importance values to each input feature.
@@ 3. Motivating Example
Fig 1(a) presents the function “parserep” in the tcpdump project, designed for parsing reply packets. This code is identified as vulnerable in Common Vulnerabilities and Exposures (CVE-2017-12898) listed in the National Vulnerability Database (NVD). The associated commit log for the fix indicates that the modification addresses bounds checking for the NFSv3 WRITE procedure. It checks not only whether the byte count is present in the captured data but also whether the length of the opaque data being written is within the captured data, furthest forward in the packet, not the item before it. Following the patch, the statement in line 13, “ND_TCHECK2(dp[0], 0);”, was replaced with “ND_TCHECK(dp[0]);”
Furthermore, Fig 1(b) exhibits the function “voutf” within the curl project, which is associated with CVE-2018-16842. The commit log for this fix clarifies that it corrects an issue involving improper arithmetic when outputting warnings to stderr. Specifically, we can observe that in line 26, “len -= cut”has been replaced with “len -= cut + 1;”.
@@
This observation may raise questions for users, leading them to question the reliability of explanation outputs produced by explanation approaches as well as deep learning-based detectors. Given the limited availability of systematic studies on explanation approaches in the context of vulnerability detection, our goal is to conduct this evaluation study to address this gap in the existing literature.
Figure 1. An illustrative example which is featured by CVE-2017-12898 and CVE-2018-16842. In CVE-2017-12898, a modification in line 13 can be observed after patching, and in CVE-2018-16842, line 26 can be seen to be modified. The Detector Reveal successfully identifies these two vulnerable functions. However, when employing five graph-based explainers to ascertain root causes, none of them highlight statement 13 in CVE-2017-12898 or statement 26 in CVE-2018-16842 in connection to the vulnerability.

### 21 Feb 2024 v2

@@ Abstract.
Abstract: Vulnerability detectors based on deep learning (DL) models have proven their effectiveness in recent years. However, the shroud of opacity surrounding the decision-making process of these detectors makes it difficult for security analysts to comprehend. To address this, various explanation approaches have been proposed to explain the predictions by highlighting important features, which have been demonstrated effective in domains such as computer vision and natural language processing. Unfortunately, there is still a lack of in-depth evaluation of vulnerability-critical features, such as fine-grained vulnerability-related code lines, learned and understood by these explanation approaches. In this study, we first evaluate the performance of ten explanation approaches for vulnerability detectors based on graph and sequence representations, measured by two quantitative metrics including fidelity and vulnerability line coverage rate. Our results show that fidelity alone is insufficient for evaluating these approaches, as fidelity incurs significant fluctuations across different datasets and detectors. We subsequently check the precision of the vulnerability-related code lines reported by the explanation approaches and find poor accuracy in this task among all of them. This can be attributed to the inefficiency of explainers in selecting important features and the presence of irrelevant artifacts learned by DL-based detectors.
@@ 1. Introduction
The ubiquity of software in today’s digital society underpins the paramount importance of vulnerability detection techniques. As we propel further into the digital age, software permeates every corner of our lives, from healthcare to finance, transportation to communication. This pervasive deployment has led to an exponential increase in the surface area for potential attacks, amplifying the demand for robust, accurate, and efficient vulnerability detection mechanisms. Current vulnerability detection techniques (Coverity, ; RATS, ; Infer, ; CSA, ; Fortify, ; Checkmarx, ; SVF, ; devign, ; reveal, ; ivdetect, ; dwk, ; SySeVR, ; vdp, ; vulcnn, ; Russel, ), however, are often characterized by significant limitations that stymie their effectiveness. A core shortcoming lies in their limited ability to pinpoint the exact code lines that instigate vulnerabilities. Existing detection tools frequently flag entire segments of code as potentially hazardous, leading to vast sections needing review. This wide-net approach not only overwhelms developers with high rates of false positives but also sows the seeds of inefficiency and can potentially delay mitigation efforts. The ability to identify the precise lines of code responsible for a vulnerability, thus, is an issue of critical importance and poses an intricate challenge in the realm of program analysis.
@@
In our study, the overarching objective is to conduct a rigorous evaluation and exmaination of the performance of these explanation techniques within the field of fine-grained vulnerability localization. We set our sights on evaluating five graph-based explanation methodologies: GNNExplainer (gnnexplainer, ), PGExplainer (pgexplainer, ), GNN-LRP (gnnlrp, ), GradCAM (grad, ), and DeepLift (deeplift, ). These techniques are examined in conjunction with four graph-based vulnerability detectors, namely DeepWuKong (dwk, ), IVDetect (ivdetect, ), Reveal (reveal, ), and Devign (devign, ). Furthermore, we venture into the investigation of five sequence-based techniques—LRP (LRP-LSTM, ), Shap (SHAP, ), Lime (LIME, ), GradInput (grad, ), and DeepLift (deeplift, )—in relation to sequence-based vulnerability detectors such as VulDeePecker (vdp, ), SySeVR (SySeVR, ), and TokenLSTM.
@@
In our evaluation process, we initially use fidelity as the metric to evaluate existing explanation techniques, a common practice in explainable AI to assess the accuracy of generated explanations. However, we find that relying solely on fidelity leads to significant fluctuations in scores across different detectors and datasets, showing inconsistent results. To address this, we introduce an additional metric called Vulnerability Line Coverage rate (LC), mirroring MSR (ContraFlow, ), which can be divided into two sub-metrics: the Vulnerability-Triggering Code Line Coverage rate (TLC) and the Vulnerability-Fixing Code Line Coverage rate (FLC). These aim to provide a more detailed understanding of how well the explanation techniques can identify and explain vulnerabilities. However, our findings show that most explanation techniques struggle to pinpoint key vulnerability-related statements accurately, as seen in both TLC and FLC measurements (RQ2, § 5.4). For instance, when applying GNNExplainer to IVDetect on the SARD dataset, the results show a wide range of TLC values from 7% to 64%, and FLC values from 18% to 66%, despite the fidelity scores remaining relatively high between 74% and 92%. This highlights that existing approaches, even with high fidelity scores, fail to learn and reflect the critical features inherent to the vulnerabilities, indicating that a broader evaluation beyond fidelity is necessary to truly understand the effectiveness of explanation techniques in addressing vulnerabilities.
@@
Throughout our study, we have gained two in-depth understandings regarding the DL-based vulnerability detectors and explanation approaches, respectively. On the one hand, the existing explanation approaches are prone to pick up vulnerability-irrelevant features as explanations. On the other hand, DL-based detectors indeed perform vulnerability detection by diligently learning the distinctions between vulnerable and normal samples. The varying sensitivities of model-related explainers on retrained and previous models to vulnerability-fixing lines suggest that they may primarily identify feature discrepancies between positive and negative samples, rather than understanding how vulnerabilities are triggered (RQ 3.1, § 5.5.1). To further validate our point and evaluate their comprehension of vulnerabilities, we manually patch a subset of vulnerabilities, using the originally patched samples as a reference. By comparing detection accuracy between non-vulnerability samples, we anticipate minimal accuracy differences if the detectors truly understand vulnerability nuances (RQ 3.2, § 5.5.2). Surprisingly, a significant portion of manually fixed codes is misclassified as vulnerable, with fewer false alarms in the originally patched samples. Moreover, we identify a kind of irrelevant text features that notably influence the model’s predictions, collectively indicating that current detectors have not fully understood vulnerability semantics.
@@
  • Enhancing the evaluation metrics for explanation approaches. We employ the quantitative metric of fidelity and introduce additional metrics, namely TLC and FLC, to evaluate the efficacy of the explanation approaches in accurately localizing vulnerabilities. Our investigation reveals that none of the explanation approaches effectively localize vulnerabilities even when they achieve high fidelity scores.
  • Proposing a new evaluation metric for DL-based detectors. To complement the existing evaluation metrics, we introduce a new metric VUR that is focused on the impact of vulnerability-unrelated perturbations on DL-based detectors. The metric is employed to determine the proportion of perturbations unrelated to vulnerabilities. We observe that all detectors are affected to varying extents by vulnerability-unrelated perturbations, potentially causing a flip in their prediction outcomes.
  • Revealing the limitations in DL-based detectors. As deep-learning vulnerability detectors aim to identify vulnerabilities by learning distinctions between vulnerable and normal samples, we find that the models can still be misdirected by irrelevant disturbances. Furthermore, these models are unable to recognize vulnerability-fixing patterns not present in the training dataset. We also identify a specific text feature that can trigger false positives, indicating DL-based detectors primarily distinguish between vulnerable and normal samples without a full comprehension of the mechanisms underlying vulnerabilities.
@@ 2. Background
In this section, we primarily introduce the DL-based detectors and explanation approaches used in this study.
@@ 2.1. DL-based Vulnerability Detectors
In the realm of vulnerability detection, DL-based approaches have emerged as potent tools for identifying security vulnerabilities within software programs. These methodologies typically entail segmenting source code into discrete code fragments, subsequently vectorizing each fragment using deep learning models, and predicting the likelihood of a vulnerability existing within that fragment. The inherent advantage of these methods lies in their capacity to autonomously discern patterns and regularities within source code, thereby enhancing the accuracy and efficiency of vulnerability detection.
Based on how source code is segmented, DL-based vulnerability detectors can be categorized into two primary classes: function-level and slice-level detectors. In function-level detectors (Russel, ; funded, ; reveal, ; ivdetect, ; devign, ; linevd, ; linevul, ), individual functions within the program are treated as isolated code fragments. In contrast, slice-level detectors (vulspg, ; dwk, ; vdp, ; SySeVR, ; vuldeelocator, ) extract code fragments spanning multiple functions based on data dependencies or control flow dependencies of vulnerability-critical statements. This approach offers a comprehensive contextual view of vulnerabilities, thus contributing to improved detection precision.
Furthermore, concerning the vectorization of code fragments, these vulnerability detectors can be further classified as sequence-based and graph-based detectors. Sequence-based methods directly parse source code fragments into token sequences and employ sequence neural networks for vectorization. Conversely, graph-based approaches typically parse the structures of code fragments, such as abstract syntax trees (AST), control flow graphs (CFG), or program dependency graphs (PDG), and further process these graph structures using graph neural networks. This facilitates the capture of structural information and dependency relationships within the source code, thereby enhancing the effectiveness of vulnerability detection.
@@
Table 1. Overview of seven DL-based detectors.
@@ 2.2. Explanation Approaches
  • Gradient-based (grad, ) explanation methods quantify the impact of a feature, denoted as $x_{i}$, in the input sample on the model’s output, $o$, by calculating the partial derivative of $o$ concerning feature $i$. In this research, we opt for $grad_{input}$, which represents a gradient-based method and is calculated as the dot product of the input vector and the partial gradient.
  • LRP (i.e. layer-wise relevance propagation) represents another significant approach for leveraging the internal relationships within neural networks to estimate feature importance scores. In essence, LRP utilizes the backpropagation mechanism of neural networks to propagate the output scores (relevance) from the model’s output layer all the way back to the input layer. Consequently, each input feature receives a relevance value, serving as an importance score. To apply the LRP method to LSTMs and GRUs, we employ the propagation rule introduced by Arras et al (LRP-LSTM, ).
@@
Table 2. Basic information of explanation approaches.
Explanation Method Representation Introduction GNNExplainer (gnnexplainer, ) Graph GNNExplainer trains mask for edges and node features to identify meaningful input information. These masks are updated to maximize the mutual information between original and updated predictions. PGExplainer (pgexplainer, ) Graph PGExplainer trains a parameterized mask predictor to predict the edge mask value. The predictor takes embedding vectors of two nodes connected by an edge as input and produces a corresponding mask value for the edge. GNN-LRP (gnnlrp, ) Graph GNN-LRP explores the importance of various graph walks through score decomposition. It decomposes model predictions into different graph walks using LRP GradCAM (grad, ) Graph Pope et al. (grad, ) extend GradCAM (gradcam, ) to the realm of graph classification, where it initially calculates the gradients of the target prediction with respect to the final node embeddings. Subsequently, these gradients are averaged to derive the weights for each feature map. DeepLift (deeplift, ) Both DeepLift combines LRP and Integrate Gradient, utilizing their insights to enhance its backpropagation process. By incorporating a reference input, DeepLift overcomes discontinuities and achieves superior performance in various tasks compared to traditional LRP methods. It is extended to support deep graph models by Yuan et al. (explain_survey, ). Lime (LIME, ) Sequence Lime employ a surrogate model to approximate the performance of the target model. By perturbing the input $x$, LIME samples neighboring points ${x_{1}^{{}^{\prime}},x_{2}^{{}^{\prime}},...,x_{n}^{{}^{\prime}}}$ and obtains their labels through the target model. Shap (SHAP, ) Sequence Shap estimates the Shapley values of each feature in the prediction by adjusting the loss function, weight kernel, and solving a regression problem, thus providing additive explanations. Gradient * Input (grad, ) Sequence Gradient * Input is an early-stage gradient-based attribution method that computes attribution by taking the partial derivative of the output with respect to the input and multiplying it by the input itself. LRP (LRP-LSTM, ) Sequence LRP is a technique that approximates feature importance scores by leveraging the internal relationship between input and output in neural networks. It uses backpropagation to propagate relevance scores from the output layer to the input layer, assigning importance values to each input feature.
@@ 3. Motivating Example
Fig. 1 (a) presents the function “parserep” in the tcpdump project, designed for parsing reply packets. This code is identified as vulnerable in Common Vulnerabilities and Exposures (CVE-2017-12898) listed in the National Vulnerability Database (NVD). The associated commit log for the fix indicates that the modification addresses bounds checking for the NFSv3 WRITE procedure. It checks whether the byte count is present in the captured data and whether the length of the opaque data being written is within the captured data, furthest forward in the packet, not the item before it. Following the patch, the statement in line 13, “ND_TCHECK2(dp[0], 0);”, was replaced with “ND_TCHECK(dp[0]);”
Furthermore, Fig. 1(b) exhibits the function “voutf” within the curl project, which is associated with CVE-2018-16842. The commit log for this fix clarifies that it corrects an issue involving improper arithmetic when outputting warnings to stderr. Specifically, we can observe that in line 26, “len -= cut”has been replaced with “len -= cut + 1;”.
@@
This observation may raise questions for users, leading them to question the reliability of explanation outputs produced by explanation approaches as well as DL-based detectors. Given the limited availability of systematic studies on explanation approaches in the context of vulnerability detection, our goal is to conduct this evaluation study to address this gap in the existing literature.
Figure 1. An illustrative example featured by CVE-2017-12898 and CVE-2018-16842. In CVE-2017-12898, a modification in line 13 can be observed after patching, and in CVE-2018-16842, line 26 can be seen to be modified. The detector Reveal successfully identifies these two vulnerable functions. However, when employing five graph-based explainers to ascertain root causes, none of them highlight statement 13 in CVE-2017-12898 or statement 26 in CVE-2018-16842 in connection to the vulnerability.